# Scatterless interference: Delay of laminar-to-turbulent flow transition by a lattice of subsurface phonons

### v3

@@ Abstract
Wave interference has historically relied on scattering objects placed within the wave domain. Here, we introduce a fundamentally new mechanism: scatterless interference induced by a lattice of subsurface phonon motion beneath a smooth wall interfacing with a transitioning boundary-layer flow. The subsurface consists of a wall-parallel lattice of wall-normal frequency-dependent phononic structural units, each designed to respond to local flow perturbations in an out-of-phase manner, suppressing them at the point of interaction. Collectively, the lattice induces interference effects that cause the kinetic energy of flow instabilities to decay downstream, thereby delaying laminar-to-turbulent transition. To guide the design of the phononic subsurface lattice, a Bloch-wave unit-cell analysis is developed for the flow perturbations, and direct numerical simulations validate the concept. This work establishes scatterless interference as a distinct physical phenomenon and represents a paradigm shift in the design of aerodynamic and hydrodynamic surfaces—moving beyond streamlined shaping to leveraging subsurface phonon engineering for drag reduction and enhanced performance.
@@ Introduction
The control of wave interference through the deliberate placement of scattering objects has been a cornerstone principle in wave physics for over two centuries. Its classical origins trace back to Thomas Young’s double-slit experiment [1], which first demonstrated the wave nature of light via interference between scattered paths. This concept evolved into a general principle applied across multiple disciplines. In electromagnetics, the development of photonic crystals [2, 3] relied on arrays of dielectric scatterers directly interacting with the wave field to trigger spatial constructive and destructive interferences that bring rise to Bragg scattering. A similar mechanism emerged in acoustics and elasticity, where periodic voids or inclusions, forming phononic crystals [4, 5], have enabled the formation of wave interference patterns and spectral band gaps through in-domain scattering. Even in quantum mechanics, interference effects typically arise through the interaction of wavefunctions with potentials or obstacles embedded in the same domain of motion. These classical and modern mechanisms share a foundational attribute: the structures responsible for interference are located within the same spatial domain as the wave field they manipulate  [^1] The only exception is having active external excitations to trigger the scattering [55, 56, 57]. To date, this paradigm has remained essentially unchallenged across physics. In this work, we introduce a fundamentally new mechanism of passive wave interference that breaks from the conventional paradigm: scatterless interference induced by a lattice of phononic subsurface units. In sharp contrast to prevalent scattering mechanisms, the interference patterns in our fully passive system arise not from scatterers embedded within the wave domain, but from elastic phononic structures buried beneath the surface—outside the spatial domain occupied by the propagating waves. We present this new physical mechanism in the context of the long-standing classical fluid dynamics problem of laminar-to-turbulent transition.
@@
Numerous research approaches have been proposed in the last couple of decades aimed at passively mitigating the undesirable effects of fluid-structure interactions. These include installation of riblets over the surface [14, 15], creation of artificial surface roughness [16, 17], drilling of holes to form a porous surface [18, 19], or coating the surface with a compliant (low-stiffness) material [20, 21, 22, 23, 24, 25, 26]. These approaches, however, are not based on a synchronization with the frequency, phase, and wavevector characteristics of the flow instabilities, and are therefore limited in their effectiveness. An ideal intervention requires a tailored solution, with mechanistic precision, to create a passive and responsive control stimulus that accounts for the dynamical properties of the underlying flow transition mechanisms. In 2015, the general concept of flow control by subsurface phonons was introduced as an approach capable of achieving this level of precise wave-synchronized control of flow instabilities [27]. A phononic subsurface (PSub) comprises a synthetically designed architected material affixed beneath  [^3] Installation of a PSub may effectively be concealed because it is placed underneath the surface. In other words, an external observer viewing a PSub-installed airplane wing would view a normally looking surface and would not be able to identify if a PSub is present or not. the surface exposed to the flow (e.g., of a wing or vehicle body). The function of the PSub is to manipulate small-amplitude vibrations on the surface, and by extension the flow perturbations near the wall that are responsible for transition. PSubs may be designed to passively respond to the flow instabilities in an out-of-phase manner, creating stabilization, or in an in-phase manner, creating destabilization--either function is realized a priori by design, or in the future may be switched or tuned in real time by actuation. In past computational investigations, a PSub has been applied as a solitary unit [27, 29, 30, 31, 32] or as a contiguous layout of units distributed along the streamwise direction [33, 34].
@@ Results
Figure 3: (a) Dispersion curves for selected unstable flow mode in the presence of a PSub lattice. The first and second rows correspond to the imaginary and real components of the wavenumber $\kappa$ for the rigid-wall case and the PSub cases in the full-span, square-lattice, and hexagonal-lattice configurations [depicted by the (b), (c) and (d) unit cell schematics, respectively]. The third row shows the real and imaginary components of the PSub’s admittance $Y$ at different frequencies. The dashed vertical lines indicate the frequencies for which the average modal perturbation KE production rate, $P_{\text{r}}$, distributions on the right plots have been obtained. Results for each frequency are depicted in their corresponding panel. The first panel (in orange) shows a weak destabilizing case (at a frequency of 1539.5 Hz) and the second panel (in light green) is for a weak stabilizing case (at a frequency of 1555 Hz). The curves below the panels correspond to the averaged production rate evolution obtained for both the weak destabilizing and weak stabilizing cases compared to that of the rigid case. As a reference, the average has been evaluated in the vicinity of the PSub wall (in the range of 10% of half of the channel’s height), and all the modal amplitudes have been normalized to make them coincide with that of the rigid-wall case at the left edge of the first unit cell.
@@ Discussion
In conclusion, the mechanism of scatterless interference reported here reveals a unique class of wave behavior that to our knowledge has not been previously reported in any branch of physics. The concept comprises subsurface phononic units arranged with full spatial flexibility and designed to vibrate at specific frequencies, enabling precise, tunable, and frequency-selective interference patterns in the overlying wave field. The physical platform we employ to demonstrate this mechanism is an unstable laminar flow field subject to perturbation waves—a canonical setting for studying aerodynamic or hydrodynamic instability. We have shown that by engineering the phonon properties of the subsurface--virtually entirely concealed beneath the flow-exposed surface--it is possible to exert full command over the very nature of the underlying fluid-structure interaction. This passive intervention offers precise mechanistic and tunable control of the behavior of wall-bounded flows in general. Furthermore, our findings demonstrate that the principles of Bloch wave analysis, traditionally applied to electronic, photonic, acoustic, and elastic waves [36, 37], can be effectively extended to flow perturbations influenced by PSubs. By leveraging locally resonant metamaterial-based individual PSub units, we have demonstrated that flow instabilities can be manipulated through interference effects, in a manner analogous to wave phenomena in periodic elastic or acoustic media. This approach enables passive control over critical processes such as the transition to turbulence, with the PSub lattice facilitating attenuation of unstable T-S waves downstream, while being robust to variations in wavenumber and direction of propagation. Other types of flow instabilities may be similarly treated in future studies. With the ongoing influx of new concepts from phonon engineering and metamaterials [38, 39, 40], the prospects of further future improvements of the performance following this framework is highly accessible, and so is the path to practical implementation given the rapid advancements in fabrication technology [41, 42]. A wide range of flow-control applications stand to benefit, spanning from hypersonic airflows to low-speed liquid transport. It is conceivable that within this framework future aircrafts will maintain a fully laminar boundary layer over their surfaces, enabling transformative improvements in fuel efficiency. In addition, PSub lattices may be simultaneously designed to inhibit aeroacoustic emissions, offering the additional rewards of improved passenger comfort and reduced environmental noise pollution. Similar advances will materialize for hydroacoustic emissions as well.

### v4

@@ Abstract
Wave interference has historically relied on scattering objects placed within the wave domain. Here, we introduce a fundamentally new mechanism: scatterless interference induced by a lattice of subsurface phonon motion beneath a smooth wall interfacing with an unstable laminal channel or boundary-layer flow. The subsurface consists of a wall-parallel lattice of wall-normal frequency-dependent phononic structural units, each designed to locally respond to a growing flow perturbation in an out-of-phase manner, dynamically influencing it at the point of interaction. Collectively, the lattice induces an interference effect that causes the kinetic energy of the flow instability to decay downstream, thereby delaying laminar-to-turbulent transition. To guide the design of the phononic subsurface lattice, a Bloch-wave unit-cell analysis is developed for flow perturbations, and direct numerical simulations validate the concept. This work establishes scatterless interference as a distinct physical phenomenon and represents a paradigm shift in the design of aerodynamic and hydrodynamic surfaces—moving beyond streamlined shaping to leveraging subsurface phonon engineering for drag reduction and enhanced performance.
@@ Introduction
The control of wave interference through the deliberate placement of scattering objects has been a cornerstone principle in wave physics for over two centuries. Its classical origins trace back to Thomas Young’s double-slit experiment [1], which first demonstrated the wave nature of light via interference between scattered paths. This concept evolved into a general principle applied across multiple disciplines. In electromagnetics, the development of photonic crystals [2, 3] relied on arrays of dielectric scatterers directly interacting with the wave field to trigger spatial constructive and destructive interferences that bring rise to Bragg scattering. A similar mechanism emerged in acoustics and elasticity, where periodic voids or inclusions, forming phononic crystals [4, 5], have enabled the formation of wave interference patterns and spectral band gaps through in-domain scattering. Even in quantum mechanics, interference effects typically arise through the interaction of wavefunctions with potentials or obstacles embedded in the same domain of motion. These classical and modern mechanisms share a foundational attribute: the structures responsible for interference are located within the same spatial domain as the wave field they manipulate  [^1] The only exception is having active external excitations to trigger the scattering [55, 56, 57]. To date, this paradigm has remained essentially unchallenged across physics. In this work, we introduce a fundamentally new mechanism of passive wave interference that breaks from the conventional framework: scatterless interference induced by a lattice of phononic subsurface units. In sharp contrast to prevalent scattering mechanisms, the interference patterns in our fully passive system arise not from scatterers embedded within the wave domain, but from elastic phononic structures buried beneath the surface—outside the spatial domain occupied by the propagating waves. We present this new physical mechanism in the context of the long-standing classical fluid dynamics problem of laminar-to-turbulent transition.
@@
Aside from the conventional practice of streamline shaping of surfaces, numerous research approaches have been proposed in the last couple of decades aimed at passively mitigating the undesirable effects of fluid-structure interactions. These include installation of riblets over the surface [14, 15], creation of artificial surface roughness [16, 17], drilling of holes to form a porous surface [18, 19], or coating the surface with a compliant (low-stiffness) viscoelestic material [20, 21, 22, 23, 24, 25, 26]. These approaches, however, are not based on a synchronization with the frequency, phase, and wavevector characteristics of the flow instabilities, and are therefore limited in their effectiveness. An ideal intervention requires a tailored solution, with mechanistic precision, to create a passive and responsive control stimulus that accounts for the dynamical properties of the underlying flow transition mechanisms. In 2015, the general concept of flow control by subsurface phonons was introduced as an approach capable of achieving this level of precise wave-synchronized control of flow instabilities [27]. A phononic subsurface (PSub) comprises a synthetically designed architected material affixed beneath  [^3] Installation of a PSub may effectively be concealed because it is placed underneath the surface. In other words, an external observer viewing a PSub-installed airplane wing would view a normally looking surface and would not be able to identify if a PSub is present or not. the surface exposed to the flow (e.g., of a wing or vehicle body). The function of the PSub is to manipulate small-amplitude vibrations on the surface, and by extension the flow perturbations near the wall that are responsible for transition. PSubs may be designed to passively respond to the flow instabilities in an out-of-phase manner, creating stabilization, or in an in-phase manner, creating destabilization--either function is realized a priori by design, or in the future may be switched or tuned in real time by actuation. In past computational investigations, a PSub has been applied as a solitary unit [27, 29, 30, 31, 32] or as a contiguous layout of units distributed along the streamwise direction [33, 34].
@@ Results
Figure 3: (a) Bloch analysus or flow perturbations: Dispersion curves for selected unstable flow mode in the presence of a PSub lattice. The first and second rows correspond to the imaginary and real components of the wavenumber $\kappa$ for the rigid-wall case and the PSub cases in the full-span, square-lattice, and hexagonal-lattice configurations [depicted by the (b), (c) and (d) unit cell schematics, respectively]. The third row shows the real and imaginary components of the PSub’s admittance $Y$ at different frequencies. The dashed vertical lines indicate the frequencies for which the average modal perturbation KE production rate, $P_{\text{r}}$, distributions on the right plots have been obtained. Results for each frequency are depicted in their corresponding panel. The first panel (in orange) shows a weak destabilizing case (at a frequency of 1539.5 Hz) and the second panel (in light green) is for a weak stabilizing case (at a frequency of 1555 Hz). The curves below the panels correspond to the averaged production rate evolution obtained for both the weak destabilizing and weak stabilizing cases compared to that of the rigid case. As a reference, the average has been evaluated in the vicinity of the PSub wall (in the range of 10% of half of the channel’s height), and all the modal amplitudes have been normalized to make them coincide with that of the rigid-wall case at the left edge of the first unit cell.
@@ Discussion
In conclusion, the mechanism of scatterless interference reported here reveals a unique class of wave behavior that to our knowledge has not been previously reported in any branch of physics. The concept comprises subsurface phononic units arranged with full spatial flexibility and designed to vibrate at specific frequencies, enabling precise, tunable, and frequency-selective interference patterns in the overlying wave field. The physical platform we employ to demonstrate this mechanism is an unstable laminar flow field subject to perturbation waves—a canonical setting for studying aerodynamic or hydrodynamic instability. We have shown that by engineering the phonon properties of the subsurface--virtually entirely concealed beneath the flow-exposed surface--it is possible to exert full command over the very nature of the underlying fluid-structure interaction. This passive intervention offers precise mechanistic and tunable control of the behavior of wall-bounded flows in general. Furthermore, our findings demonstrate that the principles of Bloch wave analysis, traditionally applied to electronic, photonic, acoustic, and elastic waves [36, 37], can be effectively extended to flow perturbations influenced by PSubs. By leveraging locally resonant metamaterial-based individual PSub units—each acting in the wall-normal direction and collectively working in tandem as a subsurface lattice along the wall-parallel plan—we have demonstrated that flow instabilities can be manipulated through interference effects in a manner analogous to wave phenomena in periodic acoustic or elastic media, and more broadly for elecrtons and photons. This approach enables passive control over critical processes such as the transition to turbulence, with the PSub lattice facilitating attenuation of unstable T-S waves downstream, while being robust to variations in wavenumber and direction of propagation. Other types of flow instabilities may be similarly treated in future studies. With the ongoing influx of new concepts from phonon engineering and metamaterials [38, 39, 40], the prospects of further future improvements of the performance following this framework is highly accessible, and so is the path to practical implementation given the rapid advancements in fabrication technology [41, 42]. A wide range of flow-control applications stand to benefit, spanning from hypersonic airflows to low-speed liquid transport. It is conceivable that within this framework future aircrafts will maintain a fully laminar boundary layer over their surfaces, enabling transformative improvements in fuel efficiency. In addition, PSub lattices may be simultaneously designed to inhibit aeroacoustic emissions, offering the additional rewards of improved passenger comfort and reduced environmental noise pollution. Similar advances will materialize for hydroacoustic emissions as well. Lastly, extensions of the scatterless interference principle to other branches of wave physics such as quantum dynamics and electromagnetics may be explored.